OpenStudy (anonymous):

what are the values of x that are solutions to f(x)=0 where f(x)=-5x^2+4x+9

OpenStudy (vishweshshrimali5):

\[\large{-5x^2 + 4x + 9 =0}\] \[\large{\implies 5x^2 - 4x - 9 = 0}\]

OpenStudy (vishweshshrimali5):

Any doubt till this step ?

OpenStudy (vishweshshrimali5):

I just multiplied (-1) on both sides

OpenStudy (vishweshshrimali5):

\[\large{5x^2 +5x - 9x - 9 = 0}\]

OpenStudy (anonymous):

the choices are a -0.56,0.44 b-5,4 c-0.80,-1.80 d-1.00,-1.80

OpenStudy (vishweshshrimali5):

We will get to that later :) First tell me do you have any doubt in the steps I wrote above ?

OpenStudy (anonymous):

yes because none of those are any of the available answers

OpenStudy (vishweshshrimali5):

No no no ... I have not calculated the answer yet.. We are just doing some work to get the answer :)

OpenStudy (anonymous):

ok got ya

OpenStudy (vishweshshrimali5):

Good lets see: \[\large{5x^2 + 5x - 9x - 9 = 0}\] \[\large{\implies 5x(x +1) - 9(x+1) = 0}\]

OpenStudy (anonymous):

you asking me to distribute this?

OpenStudy (vishweshshrimali5):

\[\large{\implies (5x-9)(x+1) = 0}\] \[\large{\implies (5x-9) = 0 ~\rm{or}~(x+1) = 0}\]

OpenStudy (vishweshshrimali5):

Nope not distribution, collecting the terms

OpenStudy (anonymous):

different problem

OpenStudy (vishweshshrimali5):

Remember this: \[\large{xy = 0 \implies x = 0 ~or~y = 0}\]

OpenStudy (vishweshshrimali5):

Thus: \[\large{(5x-9)(x+1) = 0}\] \[\large{\implies 5x - 9 = 0 ~or~x+1 = 0}\]

OpenStudy (vishweshshrimali5):

Now solve each expression one by one: 1. 5x - 9 = 0 => 5x = 9 => x = 9/5 2. x+1 = 0 => x = -1

OpenStudy (vishweshshrimali5):

Any doubt ?